OpenStudy (anonymous):

hi. can someone help me explain how to get a general soln for (x^3+(y/x)dx + (y^2+lnx)dy=0?

OpenStudy (anonymous):

sorry cant use this.

OpenStudy (anonymous):

\[\left(x^3 + \frac{y}{x}\right)\:dx + (y^2 + \ln x)\:dy = 0\]my first guess would be exact DE

OpenStudy (anonymous):

Oh hey! welcome back!

OpenStudy (anonymous):

hey, exraven, could you verify the general soln in the previous thread that we were both in? I want to make sure that's correct.

OpenStudy (anonymous):

I have already written the general soln there

OpenStudy (anonymous):

oh. I plugged in the value for z like you said. did i end up with the correct soln?

OpenStudy (anonymous):

disregard that last comment. just saw that post just now.

OpenStudy (anonymous):

im going to go ahead and try the exact method and ill check back with my answer. i hope you don't mind.

OpenStudy (anonymous):

ok, go ahead

OpenStudy (anonymous):

the exact method for the new problem*

OpenStudy (anonymous):

ok, im back

OpenStudy (anonymous):

should the general soln be ln|x|+ln|x|+c

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

i said dm/dy = 1/x =dn/dx

OpenStudy (anonymous):

and then I said this proved that this is an exact eq and that there exists f(x,y) such that df/dx=M(x,y) and df/dy=N(x,y)

OpenStudy (anonymous):

and then I proceeded to integrate

OpenStudy (anonymous):

\[M = x^3 + \frac{y}{x}\]\[N = y^2 + \ln x\]\[\frac{\partial M}{\partial y} = \frac{1}{x}\]\[\frac{\partial N}{\partial x}=\frac{1}{x}\]the DE is exact

OpenStudy (anonymous):

wait! i just noticed something. the integral of df/dx with respect to x of (1/x) is -1/x^2 not lnx|

OpenStudy (anonymous):

yup. so far so good. I have what you wrote.

OpenStudy (anonymous):

then there exist the solution\[f(x,y) = C\]such that\[\frac{\partial f}{\partial x} = M\]and\[\frac{\partial f}{\partial y} = N\]

OpenStudy (anonymous):

\[\frac{\partial f}{\partial x} = x^3 + \frac{y}{x}\]integrate both sides w.r.t x\[f(x,y) = \frac{1}{4}x^4 + y \ln x + h(y)\]

OpenStudy (anonymous):

what about df/dy? i integrated that w.r.t. y and i got y^3/3 +ylnx. Also, how do we solve for h? would that eventually turn into C?

OpenStudy (anonymous):

\[h(y) = \int\limits y^2 \:dy = \frac{1}{3}y^3\]\[\frac{\partial f}{\partial y} = \ln x + h'(y) = N\]\[\ln x + h'(y) = y^2 + \ln x\]\[h'(y) = y^2\]\[h(y) = \int\limits y^2 \: dy=\frac{1}{3}y^3\]we don't need to add the constanct because the solution already has the constant C\[f(x,y) = \frac{1}{4}x^4 + y \ln x + \frac{1}{3}y^3 \]therefore, the solution\[\frac{1}{4}x^4 + y \ln x + \frac{1}{3}y^3 = C\]

OpenStudy (anonymous):

disregard the first line

OpenStudy (anonymous):

the h(y)=...? that line?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

everything is almost crystal clear. had I left that constant (ylnx) in my f(x,y), would that be wrong?